OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a blue marble, not replacing it, and picking a brown marble?

OpenStudy (sloppycanada):

\[\frac{ 7 }{ 19 }x \frac{ 1 }{ 18 }\] Solve the following fraction and simplify. Let me know what you get.

OpenStudy (anonymous):

.37x.06=.0222? (It was multiplied right?)

OpenStudy (sloppycanada):

But as a fraction what is this?

OpenStudy (sloppycanada):

Yes it was multiplied.

OpenStudy (anonymous):

according to google .0222 as a fraction is 111 / 5000 so???

OpenStudy (sloppycanada):

Eh? No not that as a fraction. 7x1 = 7 19 x 18 = 342 The fraction should be 7/342.

OpenStudy (sloppycanada):

The decimal should be .020467

OpenStudy (anonymous):

\[\frac{7}{19\ 18}=\frac{7}{342} \]